# Multi-Subarray FD-RIS Enhanced Multi-user Wireless Networks: With Joint Distance-Angle Beamforming

### v1

@@ Abstract
The concept of the frequency diverse reconfigurable intelligent surface (FD-RIS) technology has been introduced, which can enable simultaneous implementation of distance-angle beamforming in far-field communication scenarios. In order to improve the managing ability on undesired harmonic signals and the diversity of frequency offsets, this paper presents a novel multi-subarray FD-RIS framework. In this framework, the RIS is evenly divided into multiple subarrays, each employing a distinct time-modulation frequency to enable the diversity of frequency offsets. Additionally, to suppress the undesired harmonic signals, a new time-modulation technique is employed to periodically adjust the phase-shift of each element. Based on the proposed multi-subarray FD-RIS, the signal processing model is first analytically derived. To evaluate the effectiveness of the proposed multi-subarray FD-RIS, we integrate it into a multi-user communication scenario and formulate an optimization problem that aims to maximize the weighted sum rate of all users. This is achieved by jointly optimizing the active beamforming, time delays, and modulation frequencies. Subsequently, a novel iterative algorithm is proposed to effectively solve this problem with low computing complexity. Simulation results demonstrate that the proposed multi-subarray FD-RIS can significantly enhance the performance of far-field communication networks by leveraging unique distance-angle beamforming. Furthermore, to achieve same performance gains, the FD-RIS-assisted system can substantially reduce the required number of RIS elements, number of antennas, and power budget, than the conventional RIS-assisted schemes. The proposed algorithm also demonstrates a notably superiority in performance and computational complexity compared with the baseline algorithms such as semi-definite relaxation (SDR) and zero-forcing (ZF).
@@ I Introduction
The international telecommunication union (ITU) has comprehensively defined the application scenarios and performance metrics for future 6G networks in IMT-2030 [1]. Among these, a significant emphasis has been placed on the immense demand for spectral efficiency and system capacity in communication systems. However, the spectral efficiency and system capacity of existing communication systems are gradually approaching their theoretical maximums. As a result, a significant challenge for upcoming 6G communication networks is to find new ways to improve these metrics further. As the key technology of the 5G and future 6G, massive multiple-input multiple-output (MIMO) can significantly improve the spectral effeciency of communication systems [2] by equipping with a large number of antennas at the base station (BS). However, implementing massive MIMO technology in real-world applications faces several challenges, including high costs, high hardware complexity and significant energy consumption [3]. These challenges stem primarily from the need for expensive radio frequency (RF) components and advanced signal processing systems to achieve optimal performance, which poses a substantial barrier to the practical deployment of the massive MIMO in future 6G networks.
@@ I-A Related Work
Reconfigurable intelligent surface (RIS) has emerged as a pivotal technology for future 6G networks, providing a cost-effective, hardware-efficient, and energy-efficient solution to the challenges faced by massive MIMO technology. By utilizing low-cost passive elements and simple control circuitry, RIS can effectively modify the electromagnetic properties of incoming signals. Specifically, [4] and [5] provide the systematic exploration for the RIS’s potentials in enhancing energy efficiency of multi-user communication systems, conducting the research from different optimization objectives. The focus of [4] is on minimizing total energy consumption of the system while meeting all users’ quality of service (QoS) constraints, whereas [5] aims to optimize energy efficiency within a given power budget. Simulation results demonstrate that compared to traditional active relay systems, RIS significantly improves the energy efficiency of communication systems and shows superior balancing capabilities in terms of users’ QoS. Furthermore, the simulation results in [4] indicate that by achieving similar performance gains as massive MIMO systems, integrating RIS can significantly reduce the required number of antennas at BSs, thus lowering the hardware complexity and costs of the system. Subsequently, the authors in [6] explore the feasibility of the RIS to improve the coverage of mmWave networks. In particular, they derive the closed-form expression for the peak reflection power of the RIS and the downlink signal-to-interference ratio (SIR) coverage. The results demonstrate the significant advantages of RIS in enhancing coverage for mmWave communication networks. Based on the aforementioned advantages, RIS has been incorporated into various communication networks to enhance performance in domains such as secure communications [7, 8, 9], mobile edge computing (MEC) [10, 11], and integrated sensing and communications (ISAC) [12, 13].
@@
Actually, even with the assistance of RIS technology, meeting the capacity demands of future 6G far-field communications still faces significant challenges, primarily due to the following reasons: (i) Far-field communication systems must contend with substantial path loss, which is particularly pronounced at high-frequency carriers such as mmWave or terahertz bands. This severe attenuation significantly limits both communication range and system capacity. (ii) In far-field communication scenarios, the traditional RIS can only perform beamforming for incoming signals in the angular dimension due to the plane wave properties of electromagnetic waves. This single-dimensional control significantly limits the theoretical upper bound of signal enhancement, making it difficult to further improve system energy efficiency. Addressing the challenge of path loss is inherently difficult, as it is a fundamental characteristic of signal propagation and cannot be entirely avoided. Therefore, enhancing the energy efficiency of the system becomes the key to increase the capacity of far-field communication systems. Note that [14] highlights that the RIS is able to flexibly manipulate and concentrate signal energy from both distance and angle dimensions in near-filed communications due to the unique spherical wave properties of electromagnetic waves. This dual-dimensional control significantly improves spatial resolution and energy efficiency. Motivated by the beamforming capabilities of RIS-assisted near-field communications, integrating distance-based modulation into far-field communications offers significant potential benefits. However, achieving distance-based modulation in the far-field is inherently challenging due to the planar propagation characteristics of electromagnetic waves. Therefore, a critical question arises: “How can distance-based modulation be effectively incorporated into RIS-assisted far-field beamforming?” It is worth noting that addressing this challenge is essential for unlocking new possibilities in RIS-supported far-field communication systems.
@@
Fortunately, the frequency diverse array (FDA) antenna serves as a prime example of achieving effective distance-angle beamforming in far-field communications [15]. Specifically, before transmitting the signal, a slight frequency offset is introduced to the carrier of each antenna, through which the distance dimension can be effectively incorporated into the transmitted signal. Actually, FDA antennas were originally applied in radar systems to enhance their detection capabilities [16, 17]. To further improve the spectral efficiency and capacity of wireless communications, some researchers begin to introduce FDA technology into communication systems [18, 19, 20], exploring its potentials in enhancing communication performance. [18] initially integrates the FDA technology into simultaneous wireless information and power transfer systems, and extensive numerical results demonstrate that FDA can significantly outperform phased arrays in enhancing users’ achievable rates and energy harvesting capabilities. Inspired by the intrinsic mechanism of the FDA antennas, if RIS can integrate different frequency offsets into the incident signals, it has the potential to achieve joint distance-angle beamforming for the incoming signals. However, unlike FDA antennas, RIS, as a passive device, lacks the ability to attain signal frequency diversity through RF components.
@@
To address this issue, the frequency diverse RIS (FD-RIS) proposed in [21] provides a feasible solution. Its core idea can be summarized as follows: when time modulation techniques are used to periodically alter the phase shifts of RIS, the incident signal will be scattered into a series of harmonic signals according to Fourier series expansion theory. Each harmonic frequency has a frequency offset relative to the centre carrier frequency, and the size of this offset is determined by the harmonic order and the time modulation period. Based on this principle, the authors of [21] have verified the joint distance-angle beamforming capability of the FD-RIS through both theoretical analysis and numerical simulations. Moreover, the comprehensive simulations conducted in [21] demonstrate that the FD-RIS can concentrate over threefold signal energy towards users in contrast to conventional RIS, thereby significantly improve the energy efficiency.
@@ I-B Motivation and Contributions
  • Substantial Performance Improvement: Extensive simulation analysis are conducted to verify the potential applications of deploying the designed multi-subarray FD-RIS in multi-user communication scenarios and the efficacy of the proposed iterative algorithm. The simulation results indicate that compared to traditional RIS, multi-subarray FD-RIS exhibits significant advantages in enhancing system performance. Furthermore, under the same performance gains, multi-subarray FD-RIS can significantly reduce the required number of elements, the number of antennas at the BS, and the BS’s transmit power. The comparison with semi-definite relaxation (SDR) and zero-forcing (ZF) algorithms validate the superiority of the proposed algorithm in terms of computational complexity and performance enhancement.
[…]
II Signal Processing Model for Multi-subarray FD-RIS
@@
In this section, we will establish the signal processing model of the multi-subarray FD-RIS. Specifically, it is assumed that the multi-array FD-RIS is constituted by $L=R\times S$ subarrays, where each subarray possesses $M\times N$ elements and a time modulation period $T_{l}$, for $l\in\mathcal{L}\triangleq\{1,\cdots,L\}$. Hence, the total number of elements equipped at the FD-RIS is $I=I_{z}\times I_{y}$, where $I_{z}=R\times M$ and $I_{y}=S\times N$. When the BS with $N_{\mathrm{t}}$ antennas transmits the signal $\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}e^{j2\pi f_{\mathrm{c}}t}s$ to the FD-RIS, with $w_{\mathrm{n_{\mathrm{t}}}}$ representing the beamforming coefficient of the $n_{\mathrm{t}}$-th antenna, $f_{\mathrm{c}}$ denoting the centre carrier frequency, and $s$ being the narrow-band complex envelope, the FD-RIS with multiple subarrays will reflect the incident signal as
  • $x(t)=\zeta(d_{\mathrm{br}})s\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}\sum_{i_{z}=1}^{I_{z}}\sum_{i_{y}=1}^{I_{y}}\\ e^{j2\pi f_{\mathrm{c}}\left(t-\frac{d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}}{c}\right)}\Upsilon_{i_{z},i_{y}}(t),$
@@
where $\zeta(d_{\mathrm{br}})$ represents the large-scale path loss of the transmitted signal from the BS to the FD-RIS with $d_{\mathrm{br}}$ being the spatial space between the reference antenna at the BS and the reference element at the FD-RIS. $\Upsilon_{i_{z},i_{y}}(t)$ indicates the time-modulation reflected coefficient of the $(i_{z},i_{y})$-th element of the FD-RIS. $d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}$ is the distance between the $n_{\mathrm{t}}$-th antenna and the $(i_{z},i_{y})$-th element, which is given by [22] as $d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}=d_{\mathrm{br}}+\Gamma_{n_{\mathrm{t}},i_{z},i_{y}}^{\mathrm{br}}$, where $\Gamma^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}=(i_{z}-1)d\cos(\theta_{\mathrm{br}})+(i_{y}-1)d\sin(\theta_{\mathrm{br}})\cos(\varphi_{\mathrm{br}})+(n_{\mathrm{t}}-1){\color[rgb]{1.00,0.00,0.00}d}\sin(\theta_{\mathrm{br}})\sin(\varphi_{\mathrm{br}})$ with $\theta_{\mathrm{br}}$ and $\varphi_{\mathrm{br}}$ being respectively the azimuth and elevation angles of arrival (AoA) for signals from the BS to the FD-RIS.
@@
  • $x(t)=A_{0}e^{j\phi_{0}}\zeta(d_{\mathrm{br}})s\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}\sum_{i_{z}=1}^{I_{z}}\\ \sum_{i_{y}=1}^{I_{y}}e^{j2\pi(f_{\mathrm{c}}+g{f}_{l})t}e^{-\frac{j2\pi d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}}{\lambda}},$
@@
To overcome this challenge, we introduce a time delay, $\tau_{i},i\in\mathscr{I}\triangleq\{1,\cdots,I\}$, into the periodical reflection coefficient of the $i$-th element in the FD-RIS. Specifically, according to the time-delay feature of the Fourier series, the reflection coefficient of the $i$-th element with a time delay $\tau_{i}$ can be expressed as
@@
where $i_{z}=\lceil\frac{i}{I_{y}}\rceil$, $i_{y}=\operatorname{mod}(i,I_{y})$, $i\in\mathscr{I}$. The above conclusion indicates that the $i$-th element can offer an arbitrary phase-shift for the reflected harmonic signal by properly and carefully introducing a time delay. Consequently, implementing time delays on each element of the FD-RIS will significantly enhance the spatial modulation capabilities of the FD-RIS for the incident signals. In this case, the reflected signal of the FD-RIS can be further expressed as
  • $x(t)=A_{0}e^{j\phi_{0}}\zeta(d_{\mathrm{br}})s\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}\sum_{i_{z}=1}^{I_{z}}\\ \sum_{i_{y}=1}^{I_{y}}e^{j2\pi(f_{\mathrm{c}}+g{f}_{l})t}e^{-\frac{j2\pi d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}}{\lambda}}e^{-2\pi g{f}_{l}\tau_{i}}.$
@@
where $d_{i_{z},i_{y}}=\hat{d}+\Gamma_{i_{z},i_{y}}$ with $\Gamma_{i_{z},i_{y}}=(i_{z}-1)d\cos(\theta)+(i_{y}-1)d\sin(\theta)\cos(\varphi)$, and $n\sim(0,\sigma^{2})$ is the additive white Gaussian noise (AWGN) with $\sigma^{2}$ being the noise power. The energy of the received signal can be calculated as
$\left|y(t)\right|^{2}=$ $A_{0}^{2}\zeta^{2}(\hat{d})\zeta^{2}(d_{\mathrm{br}})\Bigg{|}\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}\sum_{i_{z}=1}^{I_{z}}\sum_{i_{y}=1}^{I_{y}}e^{j2\pi g{f}_{l}(t-\frac{d_{iz,iy}}{c})}$
$e^{-j2\pi g{f}_{l}\tau_{i}}e^{-\frac{j2\pi\left(\Gamma^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}+\Gamma_{i_{z},i_{y}}\right)}{\lambda}}\Bigg{|}^{2}+\sigma^{2}.$ (9)
According to the expression of the received signal energy, we can find that the presence of the frequency offsets causes the received signal energy to be influenced not only by the angles but also by the distances between the FD-RIS and the users. This means that we can precisely manipulate the signal energy concentration towards the designated position in terms of angle and distance by meticulously designing the reflection coefficients of the FD-RIS, thereby augmenting the signal-to-interference-plus-noise ratio (SINR) at the specified location.

### v2

@@ Abstract
The concept of the frequency diverse reconfigurable intelligent surface (FD-RIS) technology has been introduced, which can enable simultaneous implementation of distance-angle beamforming in far-field communication scenarios. In order to improve the managing ability on undesired harmonic signals and the diversity of frequency offsets, this paper presents a novel multi-subarray FD-RIS framework. In this framework, the RIS is evenly divided into multiple subarrays, each employing a distinct time-modulation frequency to enable the diversity of frequency offsets. Additionally, to suppress the undesired harmonic signals, a new time-modulation technique is employed to periodically adjust the phase-shift of each element. In this case, the signal processing model is first analytically derived. Then, we integrate it into a multi-user communication scenario and formulate an optimization problem that aims to maximize the weighted sum rate of all users. This is achieved by jointly optimizing the active beamforming, time delays, and modulation frequencies. Subsequently, a novel iterative algorithm is proposed to effectively solve this problem with low computing complexity. To evaluate the beamforming capability of the proposed multi-subarray FD-RIS, three communication scenarios with varying spatial correlations among users are considered. Simulation results demonstrate that the proposed multi-subarray FD-RIS can significantly enhance and sustain the performance of communication networks by leveraging unique distance-angle beamforming, even when users share the same angular position where traditional RIS experiences severe degradation. The proposed algorithm also demonstrates a notably superiority in performance and computational complexity compared with the baseline algorithms such as semi-definite relaxation (SDR) and zero-forcing (ZF).
@@ I Introduction
The international telecommunication union (ITU) has comprehensively defined the application scenarios and performance metrics for future 6G networks in IMT-2030 [1]. Among these, a significant emphasis has been placed on the immense demand for spectral efficiency and system capacity in communication systems. However, the spectral efficiency and system capacity of existing communication systems are gradually approaching their theoretical maximums. As a result, a significant challenge for upcoming 6G communication networks is to find new ways to improve these metrics further. As a key technology for 5G and future 6G networks, massive multiple-input multiple-output (MIMO) can significantly enhance the spectral efficiency of communication systems [2] by deploying a large number of antennas at the base station (BS). Leveraging these antennas, massive MIMO achieves strong beamforming capability, which in turn greatly improves energy efficiency of systems and interference management ability among users. However, implementing massive MIMO technology in real-world applications faces several challenges, including high costs, high hardware complexity and significant energy consumption [3]. These challenges stem primarily from the need for expensive radio frequency (RF) components and advanced signal processing systems to achieve optimal performance, which poses a substantial barrier to the practical deployment of the massive MIMO in future 6G networks.
@@ I-A Related Work
Reconfigurable intelligent surface (RIS) has emerged as a pivotal technology for future 6G networks, providing a cost-effective, hardware-efficient, and energy-efficient solution to mitigate the challenges faced by massive MIMO technology. By utilizing low-cost passive elements and simple control circuitry, RIS can effectively modify the electromagnetic properties of incoming signals, thereby enabling powerful beamforming capability [4, 5, 6, 7]. Specifically, [4] and [5] provide the systematic exploration for the RIS’s potentials in enhancing energy efficiency of multi-user communication systems, conducting the research from different optimization objectives. The focus of [4] is on minimizing total energy consumption of the system while meeting all users’ quality of service (QoS) constraints, whereas [5] aims to optimize energy efficiency within a given power budget. Simulation results demonstrate that compared to traditional active relay systems, RIS significantly improves the energy efficiency of communication systems and shows superior balancing capabilities in terms of users’ QoS. Furthermore, the simulation results in [4] indicate that by achieving similar performance gains as massive MIMO systems, integrating RIS can significantly reduce the required number of antennas at BSs, thus lowering the hardware complexity and costs of the system. Subsequently, the authors in [6] explore the feasibility of the RIS to improve the coverage of mmWave networks. In particular, they derive the closed-form expression for the peak reflection power of the RIS and the downlink signal-to-interference ratio (SIR) coverage. The results demonstrate the significant advantages of RIS in enhancing coverage for mmWave communication networks. Based on the aforementioned advantages, RIS has been incorporated into various communication networks to enhance performance such as secure communications [8, 9, 10], mobile edge computing (MEC) [11, 12, 13, 14], and integrated sensing and communications (ISAC) [15, 16].
@@
Actually, the strong signal control capability of RIS is significantly constrained in certain communication scenarios  [^1] Unless stated otherwise, the system studied in this paper operates under the far-field communication model., particularly when some users exhibit high spatial correlation or even overlap in spatial directions. This limitation stems from the fact that conventional RIS can perform beamforming for incident signals only in the angular domain due to the plane-wave characteristics of electromagnetic propagation, which poses significant challenges in enabling independent beamforming or differentiated signal control for such users. Therefore, the single-dimensional control of conventional RIS lacks the adaptability and robustness for multi-user scenarios. To address this issue, introducing new beamforming degrees of freedom (DoFs) represents a promising research direction. Note that [17] highlights that the RIS is able to flexibly manipulate and concentrate signal energy from both distance and angle dimensions in near-field communications due to the unique spherical wave properties of electromagnetic waves. This dual-dimensional control significantly improves spatial resolution and energy efficiency. Motivated by the beamforming capabilities, integrating distance-based beamforming into RIS holds significant potential for overcoming the aforementioned challenges. However, realizing distance-based beamforming inherently difficult due to the planar propagation characteristics of electromagnetic waves. Therefore, a critical question arises: “How can distance-based beamforming be effectively incorporated into RIS’s beamforming?” It is worth noting that addressing this challenge is essential for unlocking new possibilities in RIS-supported communication systems.
@@
Fortunately, the frequency diverse array (FDA) antenna serves as a prime example of achieving effective distance-angle beamforming in antenna systems [18]. Specifically, before transmitting the signal, a slight frequency offset is introduced to the carrier of each antenna, through which the distance-dimension beamforming can be effectively incorporated into the signal’s transmitting. Actually, FDA antennas were originally applied in radar systems to enhance their detection capabilities [19, 20]. Recently, some researchers begin to introduce FDA technology into communication systems [21, 22, 23] to explore its potentials in enhancing communication performance. [21] initially integrates the FDA technology into simultaneous wireless information and power transfer systems, and extensive numerical results demonstrate that FDA can significantly outperform phased arrays in enhancing users’ achievable rates and energy harvesting capabilities. Inspired by the intrinsic mechanism of the FDA antennas, if RIS can integrate different frequency offsets into the incident signals, it has the potential to achieve joint distance-angle beamforming for the incoming signals. However, unlike FDA antennas, RIS, as a passive device, lacks the ability to attain signal frequency diversity through RF components.
@@
To address this issue, the frequency diverse RIS (FD-RIS) proposed in [24] provides a feasible solution. Its core idea can be summarized as follows: when time modulation techniques are used to periodically alter the phase shifts of RIS, the incident signal will be scattered into a series of harmonic signals according to Fourier series expansion theory. Each harmonic frequency has a frequency offset relative to the centre carrier frequency, and the size of this offset is determined by the harmonic order and the time modulation period. Based on this principle, the authors of [24] have verified the joint distance-angle beamforming capability of the FD-RIS through both theoretical analysis and numerical simulations.
@@ I-B Motivation and Contributions
  • Substantial Performance Improvement: To quantify the beamforming capability of the proposed multi-subarray FD-RIS, three communication scenarios characterized by different spatial correlation levels among users are considered. Extensive simulations are conducted and the obtained simulation results indicate that compared to traditional RIS, the proposed multi-subarray FD-RIS can substantially enhance and effectively sustain system performance through its unique distance–angle beamforming capability, particularly when users share the same spatial angle, underscoring its substantial potential for future wireless networks. The comparison with semi-definite relaxation (SDR) and zero-forcing (ZF) algorithms validate the superiority of the proposed algorithm in terms of computational complexity and performance enhancement.
The remainder of this paper is organized as follows. Section II introduces the signal processing model of the multi-subarray FD-RIS and demonstrates its distance–angle beamforming capability. Section III presents the communication system model and formulates the optimization problem. In Section IV, an iterative algorithm is developed to jointly optimize the active beamforming, time delays, and time-modulation frequencies. Section V provides simulation results to verify both the potential of the proposed FD-RIS and the effectiveness of the developed algorithm. Finally, Section VI concludes the paper and offers some application prospects of FD-RIS.
Notation: $\lceil\cdot\rceil$ denotes the ceiling operator, which rounds a number up to the nearest integer. Operator $\circ$ signifies the Hadamard product. The operations $(\cdot)^{T}$, $(\cdot)^{*}$ and $(\cdot)^{H}$ denote the transpose, conjugate and conjugate transpose, correspondingly. $\operatorname{Diag}(\mathbf{a})$ denotes a diagonal matrix whose diagonal elements are composed of the vector $\mathbf{a}$ . $\operatorname{diag}(\mathbf{A})$ refers to a vector whose components consist of the diagonal elements of matrix $\mathbf{A}$. Additionally, the symbols $|\cdot|$ and $\|\cdot\|$ are indicative of the complex modulus and complex vector modulus, respectively. $\mathcal{I}_{i}(\mathbf{a})$ represents the $i$-th entry of the vector $\mathbf{a}$. $\mathbf{A}^{\dagger}$ represents the pseudo-inverse of the matrix $\mathbf{A}$. $x\sim\mathcal{CN}(c,d)$ represents a circularly symmetric complex Gaussian random variable denoted by $x$, characterized by a mean of $c$ and a variance of $d$. $(x)^{+}$ and $(x)^{-}$ denote the operations of $\max(0,x)$ and $\max(0,-x)$, respectively.
II Signal Processing Model
@@
In this section, we will establish the signal processing model of the multi-subarray FD-RIS with the uniform planar array (UPA) structure. Specifically, it is assumed that the multi-array FD-RIS is constituted by $L=R\times S$ subarrays, where each subarray possesses $M\times N$ elements and a time modulation period $T_{l}$, for $l\in\mathcal{L}\triangleq\{1,\cdots,L\}$. Hence, the total number of elements equipped at the FD-RIS is $I=I_{z}\times I_{y}$, where $I_{z}=R\times M$ and $I_{y}=S\times N$. When the BS with $N_{\mathrm{t}}$ antennas transmits the signal $\mathbf{x}_{\mathrm{BS}}(t)=[w_{1}e^{j2\pi f_{\mathrm{c}}t}s,\cdots,w_{n_{\mathrm{t}}}e^{j2\pi f_{\mathrm{c}}t}s,\cdots,w_{N_{\mathrm{t}}}e^{j2\pi f_{\mathrm{c}}t}s]^{T}$ to the FD-RIS, with $w_{\mathrm{n_{\mathrm{t}}}}$ representing the beamforming coefficient of the $n_{\mathrm{t}}$-th antenna, $f_{\mathrm{c}}$ denoting the centre carrier frequency, and $s$ being the narrow-band complex envelope, the signals transmitted by different antennas are linearly superimposed in space due to the linearity of the wireless channel. Hence, for each FD-RIS element, the received signal can be viewed as $\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{n_{\mathrm{t}}}e^{j2\pi f_{\mathrm{c}}t}s.$ Therefore, the signals reflected by the FD-RIS with multiple subarrays will be superimposed as
$x(t)=$ $\zeta(d_{\mathrm{br}})s\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}\sum_{i_{z}=1}^{I_{z}}\sum_{i_{y}=1}^{I_{y}}$
$e^{j2\pi f_{\mathrm{c}}\left(t-\frac{d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}}{c}\right)}\Upsilon_{i_{z},i_{y}}(t),$ (1)
@@
where $\zeta(d_{\mathrm{br}})$ represents the large-scale path loss of the transmitted signal from the BS to the FD-RIS with $d_{\mathrm{br}}$ being the spatial space between the reference antenna at the BS and the reference element at the FD-RIS. $\Upsilon_{i_{z},i_{y}}(t)$ indicates the time-modulation reflected coefficient for the $(i_{z},i_{y})$-th element of the FD-RIS. $d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}$ is the distance between the $n_{\mathrm{t}}$-th antenna and the $(i_{z},i_{y})$-th element, which is given as $d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}=d_{\mathrm{br}}+\Gamma_{n_{\mathrm{t}},i_{z},i_{y}}^{\mathrm{br}}$ [25], where $\Gamma^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}=\Gamma^{\mathrm{br}}_{i_{z},i_{y}}+(n_{\mathrm{t}}-1)d\sin(\theta_{\mathrm{br}})\sin(\varphi_{\mathrm{br}})$, $\Gamma^{\mathrm{br}}_{i_{z},i_{y}}=(i_{z}-1)d\cos(\theta_{\mathrm{br}})+(i_{y}-1)d\sin(\theta_{\mathrm{br}})\cos(\varphi_{\mathrm{br}})$, with $\theta_{\mathrm{br}}$ and $\varphi_{\mathrm{br}}$ respectively being the azimuth and elevation angles of arrival (AoA) for signals from BS to FD-RIS, and $d$ is the space between elements of FD-RIS.
@@
$x(t)=$ $A_{0}e^{j\phi_{0}}\zeta(d_{\mathrm{br}})s\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}\sum_{i_{z}=1}^{I_{z}}$
$\sum_{i_{y}=1}^{I_{y}}e^{j2\pi(f_{\mathrm{c}}+g{f}_{l})t}e^{-\frac{j2\pi d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}}{\lambda}},$ (8)
@@
To overcome this challenge, we introduce a time delay  [^2] The time delay represent a controllable temporal shift applied to the phase-control sequence of each RIS element. Such a time-shifted control process can be readily realized by adjusting the timing of the digital control signals generated by the FPGA., $\tau_{i},i\in\mathscr{I}\triangleq\{1,\cdots,I\}$, into the periodical reflection coefficient of the $i$-th element in the FD-RIS. Specifically, according to the time-delay feature of the Fourier series, the reflection coefficient of the $i$-th element with a time delay $\tau_{i}$ can be expressed as
@@
where $i_{z}=\lceil\frac{i}{I_{y}}\rceil$, $_{y}=\operatorname{mod}(i,I_{y}), ii\in\mathscr{I}$. The above conclusion indicates that the $i$-th element can offer an arbitrary phase-shift for the reflected harmonic signal by properly and carefully introducing a time delay. Consequently, implementing time delays on each element of the FD-RIS will significantly enhance the spatial modulation capabilities of the FD-RIS for the incident signals. In this case, the superimposition form of the reflected signals of the FD-RIS can be further expressed as
$x(t)=$ $A_{0}e^{j\phi_{0}}\zeta(d_{\mathrm{br}})s\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}w_{\mathrm{n_{\mathrm{t}}}}\sum_{i_{z}=1}^{I_{z}}$
$\sum_{i_{y}=1}^{I_{y}}e^{j2\pi(f_{\mathrm{c}}+g{f}_{l})t}e^{-\frac{j2\pi d^{\mathrm{br}}_{n_{\mathrm{t}},i_{z},i_{y}}}{\lambda}}e^{-j2\pi g{f}_{l}\tau_{i}}.$ (10)
@@
where $d_{i_{z},i_{y}}=\hat{d}+\Gamma_{i_{z},i_{y}}$ with $\Gamma_{i_{z},i_{y}}=(i_{z}-1)d\cos(\theta)+(i_{y}-1)d\sin(\theta)\cos(\varphi)$, and $n\sim(0,\sigma^{2})$ is the additive white Gaussian noise (AWGN) with $\sigma^{2}$ being the noise power.
Next, the distance-angle beamforming of the proposed multi-subarray FD-RIS is verified. Specifically, we consider the case with $N_{\mathrm{t}}=1$ and temporarily ignore the large-scale path loss. Based on the received signal expression in (11), the normalized beamforming gain can be derived as
$BF=$ $\frac{1}{I^{2}}\Bigg|\sum_{i_{z}=1}^{I_{z}}\sum_{i_{y}=1}^{I_{y}}e^{j2\pi g{f}_{l}(t-\frac{d_{iz,iy}}{c})}\times$
$e^{-j2\pi g{f}_{l}\tau_{i}}e^{-\frac{j2\pi\left(\Gamma^{\mathrm{br}}_{i_{z},i_{y}}+\Gamma_{i_{z},i_{y}}\right)}{\lambda}}\Bigg|^{2}.$ (12)
Then, we aim at maximizing the normalized beamforming gain $BF$ by appropriately selecting time delays, $\tau_{i},i\in\mathscr{I}$, and time-modulation frequencies $f_{l},l\in\mathcal{L}$. Fig. 1 shows the optimal normalized beampattern of FD-RIS and traditional RIS. It is observed that conventional RIS can manipulate signal beamforming only in the angular domain. However, this single-dimensional capability becomes severely limited when multiple users are located at the same angle, as the system cannot effectively distinguish between them, thereby degrading their QoS. In contrast, FD-RIS enables simultaneous control in both the distance and angular domains, offering new opportunities to address this limitation. By leveraging dual-domain control, even when users share the same angular position, the system can exploit differences in the distance domain to reshape signal propagation. With this flexible spatial manipulation, FD-RIS shows great potential to overcome the beamforming limitations of conventional RIS.
Figure 1: Optimal normalized beampattern considering $(\hat{d},\theta,\varphi)=(70m,50^{\circ},40^{\circ})$, $f_{l}=0.2+20(l-1)/(L-1)$ MHz, $l\in\mathcal{L}$: (a) Conventional RIS; (b) FD-RIS.